OpenStudy (anonymous):

Suppose a certain population increases by 20% every ten years. By what percent does the population increase each year (assuming a constant growth rate)? (Hint: the answer is not 2%.) Show your work. Round to the nearest hundredth of a percent.

OpenStudy (kropot72):

Let the annual growth rate as a decimal = x \[(1+x)^{10}=1.2\] Taking natural logs of both sides we get \[10 \times \ln (1+x)=\ln 1.2\] \[\ln (1+x)=\frac{\ln 1.2}{10}=0.018232\] \[e ^{0.018232}=1+x\ .............(1)\] Can you now solve equation (1) to find x? When you have the value of x, multiply by 100 to convert to a decimal.

OpenStudy (anonymous):

1.83992176?

OpenStudy (kropot72):

Good work! When rounded to the nearest hundredth of a percent: x = 1.84%

OpenStudy (anonymous):

Awesome! Thanks! That actually makes sense! :)

OpenStudy (kropot72):

You're welcome :)

OpenStudy (anonymous):

Can you help me with this one? I'm not very good at the log things. :/

OpenStudy (kropot72):

The expression needs to be put into the following form first; \[\ln x ^{2}-\ln y ^{6}+\ln e ^{\frac{12}{3}}\] Does that make sense?

OpenStudy (anonymous):

Yes, that makes sense.

OpenStudy (kropot72):

Good :) Now use the following properties of logs: \[\log_{} a+\log_{} b=\log_{} (a \times b)\] \[\log_{} a-\log_{} b=\log_{} \frac{a}{b}\]

OpenStudy (anonymous):

I don't think I did this right.. I got it down to log(24/18)

OpenStudy (kropot72):

The answer must include x and y, the reason being that no values are given for these. The first two terms can be reduced to one term as follows: \[\ln x ^{2}-\ln y ^{6}=\ln (\frac{x ^{2}}{y ^{6}})\] So we now have \[\ln (\frac{x ^{2}}{y ^{6}})+\ln e ^{4}\] Can you use the properties of logs to make this the log of one quantity?

OpenStudy (anonymous):

ohh. woops. okay, so \[((x ^{2} \div y ^{6}) \times e ^{4}\]

OpenStudy (anonymous):

)

OpenStudy (anonymous):

@kropot72 ?

OpenStudy (kropot72):

I am trying to answer, but the LATEX won't work for me now. You are on the right track. Express it as ln (.....) and convert e^4 to 54.598.

OpenStudy (anonymous):

ln(x^2 / y^6) 53.598 could you distribute that to be ln (53.598x^2 / 53.598y^6)?

OpenStudy (kropot72):

Not really. Do it as ln(54.598x^2/y^6)

OpenStudy (anonymous):

okay. So that's the answer then?

OpenStudy (kropot72):

Yes.

OpenStudy (anonymous):

okay, thank you!